# Properties of squeezing functions on $h$-extendible domains

### v1

@@ 5.1. The spherically tangential convergence
Moreover, [34, Lemma $4.1$] implies that
@@
where $a=\frac{1}{2}\lim_{j\to\infty}\frac{\partial^{2}H}{\partial z\partial\bar{z}}(\alpha_{j})\epsilon_{j}^{-1}\tau_{j}^{2}>0$. Therefore, passing to a subsequence if necessary, we may assume that the sequences $\Omega_{j}:=\Delta_{j}\circ\Phi_{\eta_{j}^{\prime}}(\Omega)$ and $\Delta_{j}\circ\Phi_{\eta_{j}^{\prime}}(U_{0}^{-})$ converge normally to the Siegel half-space
@@
The remainder of the proof is to estimate $\sigma_{\Omega}(\eta_{j})$. To do this, let us first define the linear transformation $\Theta$, given by
@@
Next, let us consider the sequence of biholomorphic map $f_{j}:=\Psi\circ\Theta\circ\Delta_{j}\circ\Phi_{\eta_{j}^{\prime}}\colon\Omega\to f_{j}(\Omega)=\Psi\circ\Theta(\Omega_{j})$. Since $\Theta(0,-1)=(0,-1),\Psi(0,-1)=(0,0)$, and $\Psi(z,w)\to(0,-1)$ as $\mathcal{U}_{2}\ni(z,w)\to\infty$, it follows that for a sufficiently small $\epsilon>0$, there exists $j_{0}\in\mathbb{N}_{\geq 1}$ such that
$f_{j}(\Omega\setminus U_{0})\subset B((0,-1),\epsilon/2)\quad\text{for all }j\geq j_{0}.$
Finally, one notes that $f_{j}(\Omega\cap U_{0})$ converges normally to $\mathbb{B}^{2}$ and $f_{j}(\partial\Omega\cap U_{0})$ converges to $\partial\mathbb{B}^{2}$. Moreover,
$f_{j}(\eta_{j})=\Psi\circ\Theta(0,-1/d_{0}(\eta_{j}^{\prime}))\to(0,0)\quad\text{as }j\to\infty.$
Therefore, we may assume that
$B\big{(}(0,0),1-\epsilon\big{)}\subset F_{j}(\Omega)\subset B\big{(}(0,0),1+\epsilon\big{)},\quad\forall j\geq j_{0},$
where $F_{j}(.):=f_{j}(.)-f_{j}(\eta_{j}),\;\forall j\geq j_{0}.$ Since $F_{j}(\eta_{j})=0,\;\forall j\geq j_{0}$, it follows that
$\sigma_{\Omega}(\eta_{j})\geq\frac{1-\epsilon}{1+\epsilon},\quad\forall j\geq j_{0}.$
Since $\epsilon>0$ is arbitrary, we conclude that $\lim_{j\to\infty}\sigma_{\Omega}(\eta_{j})=1$, and the proof of Theorem 1.3 is now complete. $\Box$
@@
This ensures that $\Omega_{j}:=\phi_{{\eta}_{j}}(\mathcal{G})$ converges normally to the model $\mathcal{F}:=\{(\tilde{z},\tilde{w})\in\mathbb{C}^{2}\colon\mathrm{Re}(\tilde{w})+5|\tilde{z}|^{2}<0\}$, which is biholomorphically equivalent to the unit ball $\mathbb{B}^{2}$, and $\phi_{{\eta}_{j}}(\eta_{j})=(0,-1)\in\mathcal{F}$ for all $j\geq 1$. Therefore, by following the proof of Theorem 1.3 we conclude that $\sigma_{\mathcal{G}}(\eta_{j})\to 1$ as $j\to\infty$. $\Box$
@@
This implies that $\Omega_{j}:=\phi_{{\eta}_{j}}(\Omega)$ converges normally to the model $\mathcal{H}:=\{(\tilde{z},\tilde{w})\in\mathbb{C}^{2}\colon\mathrm{Re}(\tilde{w})+31|\tilde{z}|^{2}<0\}$, which is biholomorphically equivalent to $\mathbb{B}^{2}$, and $\phi_{{\eta}_{j}}(\eta_{j})=(0,-1)\in\mathcal{H}$ for all $j\geq 1$. By arguments as in the proof of Theorem 1.3, we conclude that $\sigma_{\Omega}(\eta_{j})\to 1$ as $j\to\infty$. $\Box$
@@
Acknowledgement.
The author was supported by the Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 101.02-2021.42.

### v2

@@ 5.1. The spherically tangential convergence
Moreover, [36, Lemma $4.1$] implies that
@@
where $a=\frac{1}{2}\lim_{j\to\infty}\frac{\partial^{2}H}{\partial z\partial\bar{z}}(\alpha_{j})\epsilon_{j}^{-1}\tau_{j}^{2}>0$. Therefore, passing to a subsequence if necessary, we may assume that the sequences $\Omega_{j}:=\Delta_{j}\circ\Phi_{\eta_{j}^{\prime}}(\Omega)$ and $\Delta_{j}\circ\Phi_{\eta_{j}^{\prime}}(\Omega\cap U_{0})$ converge normally to the Siegel half-space
@@
The remainder of the proof is to estimate $\sigma_{\Omega\cap U_{0}}(\eta_{j})$. To do this, let us first define the linear transformation $\Theta$, given by
@@
Next, let us consider the sequence of biholomorphic map $f_{j}:=\Psi\circ\Theta\circ\Delta_{j}\circ\Phi_{\eta_{j}^{\prime}}\colon\Omega\to f_{j}(\Omega)=\Psi\circ\Theta(\Omega_{j})$. Notice that $f_{j}(\Omega\cap U_{0})$ and $f_{j}(\partial\Omega\cap U_{0})$ converge normally to $\mathbb{B}^{2}$ and $\partial\mathbb{B}^{2}$, respectively. Moreover, since $\Theta(0,-1)=(0,-1)$ and $\Psi(0,-1)=(0,0)$, it follows that
$f_{j}(\eta_{j})=\Psi\circ\Theta(0,-1/d_{0}(\eta_{j}^{\prime}))=\Psi(0,-1/d_{0}(\eta_{j}^{\prime}))=\left(0,\frac{1-1/d_{0}(\eta_{j}^{\prime})}{1+1/d_{0}(\eta_{j}^{\prime})}\right)\to(0,0)\quad\text{as }j\to\infty.$
Therefore, by a similar argument as in the proof of Theorem 1.2, we conclude that for a sufficiently small $\epsilon>0$, there exists $j_{0}\in\mathbb{N}_{\geq 1}$ such that
$B\big((0,0),1-\epsilon\big)\subset F_{j}(U_{0}\cap\Omega)\subset B\big((0,0),1+\epsilon\big),\,\forall j\geq j_{0},$
where $F_{j}(\cdot):=f_{j}(\cdot)-f_{j}(\eta_{j})$ for all $j\geq j_{0}$. Since $F_{j}(\eta_{j})=0,\;\forall j\geq j_{0}$, it follows that
$\sigma_{\Omega\cap U_{0}}(\eta_{j})\geq\frac{1-\epsilon}{1+\epsilon},\quad\forall j\geq j_{0}.$
Since $\epsilon>0$ is arbitrary, we conclude that $\lim_{j\to\infty}\sigma_{\Omega\cap U_{0}}(\eta_{j})=1$, and the proof of Theorem 1.3 is now complete. $\Box$
@@
This ensures that $\Omega_{j}:=\phi_{{\eta}_{j}}(\mathcal{G}\cap U_{0})$, where $U_{0}$ is a sufficiently small neighborhood of $(0,0,0)$, converges normally to the model $\mathcal{F}:=\{(\tilde{z},\tilde{w})\in\mathbb{C}^{2}\colon\mathrm{Re}(\tilde{w})+5|\tilde{z}|^{2}<0\}$, which is biholomorphically equivalent to the unit ball $\mathbb{B}^{2}$, and $\phi_{{\eta}_{j}}(\eta_{j})=(0,-1)\in\mathcal{F}$ for all $j\geq 1$. Therefore, by following the proof of Theorem 1.3 we conclude that $\sigma_{\mathcal{G}\cap U_{0}}(\eta_{j})\to 1$ as $j\to\infty$. $\Box$
@@
This implies that $\Omega_{j}:=\phi_{{\eta}_{j}}(\Omega\cap U)$, where $U$ is a sufficiently small neighborhood of $(0,0)$, converges normally to the model $\mathcal{H}:=\{(\tilde{z},\tilde{w})\in\mathbb{C}^{2}\colon\mathrm{Re}(\tilde{w})+31|\tilde{z}|^{2}<0\}$, which is biholomorphically equivalent to $\mathbb{B}^{2}$, and $\phi_{{\eta}_{j}}(\eta_{j})=(0,-1)\in\mathcal{H}$ for all $j\geq 1$. By arguments as in the proof of Theorem 1.3, we conclude that $\sigma_{\Omega\cap U}(\eta_{j})\to 1$ as $j\to\infty$. $\Box$
@@
Acknowledgments
The author was supported by the Vietnam National Foundation for Science and Technology Development (NAFOSTED) under grant number 101.02-2021.42. We would like to thank Professor Shichao Yang for providing a counterexample that helped us complete the proof of Theorem 1.1.